The largest angle of a triangle is 6 times the size of the smallest, and the middle angle is 5 times the size of the smallest. What are the three angles measure?
Answer by Theo(13342)

largest angle = 6 times the smallest.
middle angle = 5 * the smallest.

sum of the angles of a triangle is 180.

let x = the size of the smallest angle.

the size of the middle angle is 5x.

the size of the largest angle is 6x.

x + 5x + 6x = 180

combine like terms to get 12x = 180

solve for x to get x 180/12 = 15 degrees.

5x = 75 degrees.

6x = 90 degrees.

largest angle is 90 degrees.
middle angle is 75 degrees.
smallest angle is 15 degrees.